OpenStudy (samanthagreer):

Put the following equation in slope-intercept form : 3x + 12y = -4

OpenStudy (anonymous):

what are the options

OpenStudy (samanthagreer):

there rent any

OpenStudy (samanthagreer):

there arent any

OpenStudy (mathmale):

Sam: First, I'd like for you to write the GENERAL equation for "slope-intercept form" of the equation of a straight line.

OpenStudy (samanthagreer):

i dont understnad

OpenStudy (anonymous):

y=mx+b right

OpenStudy (samanthagreer):

yes

OpenStudy (mathmale):

There are 3 forms (at least) for equations of straight lines. You should know them. Saleemah is correct. Taking his "y=mx+b," please explain the meanings of m and b.

OpenStudy (anonymous):

are you in k12 if you are what unit and lesson is this

OpenStudy (anonymous):

m=slop and b is y-intercept

OpenStudy (mathmale):

Your task in this problem is to re-write the given equation 3x + 12y = -4 in slope-intercept form, which is y=mx+b. @Saleemah.A.E : Would you please let samantha answer the questions; they were for her, not for you (although your efforts to be of help are much appreciated).

OpenStudy (anonymous):

sorry

OpenStudy (mathmale):

Given 3x + 12y = -4, we want to re-write this in the form y = mx + b. How would you (Samantha) isolate y on the left side, with the other two terms on the right side?

OpenStudy (samanthagreer):

isant it 12y=3x + -4

OpenStudy (mathmale):

Please go back to 3x + 12y = -4 and subtract 3x from both sides. You're on the right track, but you currently have two sign errors.

OpenStudy (samanthagreer):